OpenStudy (anonymous):

How many amino acids could be specified if codons consisted of two nucleotides instead of three? What problem would this present for an organism making proteins using all 20 amino acids?

OpenStudy (frostbite):

You have 4 bases that can be distributed in any order more than ones, and you need 3 bases in a codon so we can say the calculation is going to be: combinations = 4*4*4 = 4^3 = 64 Try do the calculation for 2 bases in a codon instead.

OpenStudy (frostbite):

Can give a hint and say the number you are looking for is 4^2

OpenStudy (anonymous):

im still confused

OpenStudy (frostbite):

Can you say what you are confused about?

OpenStudy (anonymous):

about the whole thing

OpenStudy (frostbite):

Okay gonna break it down part by part then :) During translation we say a codon consist of 3 bases and give a specific amino acid right? There are 4 bases the codon can consist of: G, C, A and U. Now from combinatorics we know the amount of combos you can do with 4 bases in 3 spots is: 4*4*4 = 4^3 = 64. So with 4 bases as a codon we can get all amino acids. Now assume that a codon consist of 2 bases instead, the calculation then become: 4*4 = 4^2 = 16 Here we have a problem... as there are 20 amino acids and only 16 combinations, we can't create all the amino acids.

OpenStudy (anonymous):

oh ok

OpenStudy (frostbite):

Right. Hope you can get an idea to write something from this.